# UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling

### v2

@@ Abstract
Urban region profiling aims to learn a low-dimensional representation of a given urban area while preserving its characteristics, such as demographics, infrastructure, and economic activities, for urban planning and development. However, prevalent pretrained models, particularly those reliant on satellite imagery, face dual challenges. Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place. Secondly, the lack of interpretability in pretrained models limits their utility in providing transparent evidence for urban planning. In response to these issues, we devise a novel framework entitled UrbanVLP based on Vision-Language Pretraining. Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pretrained models. Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery. Rigorous experiments conducted across six urban indicator prediction tasks underscore its superior performance.
1 Introduction
Urban Region Profiling, a scholarly endeavor within the domain of GeoAI [45, 54] and smart city [85], refers to the process of transforming urban regions into a low-dimensional vector representation using deep neural networks. Rooted in the escalating global prominence of urban environments and the imperative for urban planning, this academic discipline aims to capture and encode the multifaceted attributes of urban areas, such as demographic information, economic activity, environmental factors, infrastructure, and social dynamics, into a format that can be easily analyzed and utilized by computational models [73, 74]. Given the widespread accessibility of satellite imagery on platforms such as Google Maps [42, 51, 73], coupled with its wealth of information on regional features (e.g., road networks, building density, and vegetation coverage), a predominant majority of initiatives harness satellite imagery as the foundational modality for learning urban region representations [88].
Through a comprehensive review of the current literature, we summarize mainstream methods for urban region profiling into two categories: a) Task-specific supervised learning commonly realizes urban region profiling in a fully supervised task, e.g., identifying poverty levels [21, 59, 4, 5], crop yields [80, 70, 63, 57, 53, 78], and commercial activeness [24, 51]. However, its inherent task specificity (i.e., reliant on ample labeled data), may impede the model’s capacity for broad generalization to other downstream tasks. b) Vision pretraining, also known as a type of Self-Supervised Learning [28], aims to learn general visual features from urban satellite imagery, which are subsequently fine-tuned on a specific task for enhancing performance  [29, 6, 73, 51]. The inclusion of more auxiliary data, e.g., Points of Interests (POIs) [19, 29, 82, 40], road networks [46, 47, 12] and human mobility [71, 64, 44, 39], leads to richer, more accurate, and more useful representations of urban areas. For clarity, Figure 1(a-b) depict a sketch of these two streams of approaches.
Despite the success of using satellite imagery for urban region profiling, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture). Previous approaches in Figure 1(a-b) mainly focus on single granularity, neglecting finer-grained visual clues. Nonetheless, as the saying goes, “the devil is in the details.”. By zooming into micro levels using corresponding street-view images, a more nuanced and fine-grained understanding emerges, as shown in the left hand side of Figure 1(d). Therefore, the integration and alignment of multi-granularity information, both macro-level and micro-level, remain ongoing challenges in urban region profiling that require further exploration.
@@ 1 Introduction
Figure 1: Frameworks for urban region profiling. Compared to existing literature, we present the first attempt to introduce multi-granular visual information and high-quality calibrated texts for this task.
@@
Furthermore, while urban planners rely on explainable evidence for service deployment and urban planning, the vision-pretraining-based models for urban region profiling mostly lack interpretability. They typically learn a mapping from single/multi-modal data to the representations, posing a significant challenge in comprehending the intrinsic factors influencing the prediction outcomes. In light of the recent surge in Large Language Models (LLMs) [58, 84, 55], UrbanCLIP [74] stands out as an innovative solution, which effectively integrates textual modality as auxiliary knowledge into urban region profile, as depicted in Figure 1(c). However, the text generation process of UrbanCLIP raises several critical concerns. i) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery. ii) Homogenization: The generated descriptions appear overly simplified and general, potentially leading to homogenization and inadequate differentiation. To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
@@
In this paper, we present a Vision-Language Pretraining framework (i.e., UrbanVLP) for urban region profiling. As depicted in Figure 1(d), our model elaborately integrates multi-granularity information from both satellite (macro-level) and street-view (micro-level) imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions. Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module, which utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities. To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards. Given the impracticality of generating a large number of descriptions using GPT4V [75] or Gemini [18] due to prohibitive expense, we utilize a closely comparable GPT4V [75] model called ShareGPT4V [11]. To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric called PerceptionScore, motivated by the human evaluation system.
In summary, our major contributions lie in the following aspects:
@@
  • Multi-Granularity Cross-Modal Alignment. We explore the role of two distinct visual data modalities at various semantic granularities – satellite imagery and street-view imagery. We integrate street-view data to inject fine-grained semantic information through token-level contrastive learning. By aggregating and fusing local features within a specific region with satellite imagery data, we facilitate the development of a robust and comprehensive representation of urban landscapes.
@@
  • Automatic Text Generation and Calibration. Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism based on a newly-proposed reference-free metric, ensuring the fidelity between the generated texts and the corresponding image content.
  • Empirical Evidence. We evaluate the proposed UrbanVLP on six urban indicator prediction tasks (e.g., GDP, population, and carbon emission) in diverse urban landscapes. Extensive experiments demonstrate that our UrbanVLP outperforms existing approaches by an average improvement of 3.95% on the $R^{2}$ metric, while preserving good interpretability based on generated text.
  • New Benchmark and System. We will open-source the first multi-granular vision-text benchmark ($\mathtt{CityScape}$) for urban region profiling upon paper notification, encompassing 6 downstream tasks. Moreover, we have deployed a web platform to verify the model’s practicality (see Appendix I).
2 Related work
We formally define the problem of urban region profiling. Given a satellite images $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$. The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$. The representation can then be further utilized in downstream tasks, such as inferring urban indicators $\mathbf{Y}_{g}$ for any given region $g$. We assume that during the training phase, automated additional text descriptions are available. The detailed formulation and definitions related to this task are provided in Appendix A. Next, we introduce the related work from three aspects.
Urban Region Profiling. In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility [41, 51, 73]. For example, Urban2Vec [72] integrated street-view imagery and POI data to learn neighborhood embeddings. Some contrastive learning approaches like PG-SimCLR [73] and UrbanCLIP [74] have shown success in representing urban regions through satellite images. However, the aforementioned works exclusively considered a single type of urban imagery, overlooking the potential synergy between satellite imagery and street-view imagery, which can complement each other. Recently, some efforts [7, 27, 41, 9, 33, 8, 27] have begun to explore the simple combination of satellite imagery and street-view imagery. Vision-LSTM [27] integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification. Li et al. [41] propose utilizing street segments for region modeling and predicting socio-economic indicators by merging street-view and satellite images through a multi-stage pipeline. Nevertheless, prior studies mostly overlook distinct features in satellite and street-view images for varied granularities. To this end, we propose a unified multi-granularity cross-modal alignment framework for comprehensive urban representation in this study.
Vision-Language Pretraining (VLP). VLP aims to jointly encode vision and language in a fusion model. Early works on VLP can be broadly categorized into single-stream [13, 31, 35] and two-stream [66, 52] methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities. Milestone work CLIP [61] and its variants [43, 76] highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval. Recent works [67, 2] shift towards leveraging LLMs knowledge for vision-language representation learning. In urban region profiling, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information. Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
Large Multimodal Model (LMM). As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP [61] and ALIGN [30]. Early studies [36, 37] enhanced CLIP with refined data strategies for more diverse datasets, which have proven effective for fundamental visual tasks [38, 50, 81] but have shown limitations for complex tasks such as visual question answering. Recently, new models like MiniGPT-4 [10], LLaVA [49], and InstructBLIP [14] have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs). In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
@@ 2 Related work
3 Methodology
@@
We present an overview of our proposed framework in Figure 2, which is decoupled into two stages. In the pretraining stage, we devise an automatic text generation and calibration module using ShareGPT4V [11] to generate textual descriptions for street-view images with geographical and visual prompts. Introducing the PerceptionScore to ensure high-quality text, we then develop a multi-granularity cross-modal alignment framework through multi-level contrastive learning and fine-grained information injection. In the fine-tuning stage, we use frozen encoders from Stage 1 to extract data features, fine-tune a lightweight MLP for precise predictions, and incorporate optional street-view imagery and geographical coordinates to enrich the analysis.
3.1 Automatic Text Generation and Calibration
@@
Text Generation. For each street-view image, we utilize advanced Large Multimodal Models (LMMs) due to their robust cross-modal knowledge storage capabilities, to provide comprehensive textual descriptions. However, this confronts two primary challenges. Firstly, due to substantial financial costs associated with API usage, employing closed source LMMs like GPT4V [75] or Gemini [18] for generating tens of thousands of text descriptions is impractical. Secondly, street-view images often encompass diverse detailed information, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
@@
To address these issues, we first utilize a closely related GPT4V model, known as ShareGPT4V [11], to generate street-view image descriptions. This model is a powerful language model with vision capabilities trained on 100k GPT4V generated captions, demonstrating captioning capabilities comparable to GPT4V. In the process of designing text prompt templates, the proportion of segmented elements in street-view images reflects the distribution of elements, aiding LMMs in evaluating the importance of each element. To achieve this, we employ a pretrained segmentation model to decouple visual elements and calculate the proportion of segmentation for each element. Simultaneously, urban and geospatial coordinates of street-view images are included as prompts to guide the expression of additional urban information. Template and segmentation ratio details can be found in Appendix F.
@@
Text Calibration. As existing works [32, 16, 74] demonstrated that the quality of the textural descriptions is crucial for enhancing model performance. However, previous research [74] primarily employed simplistic rules and manual refinement processes for text description of satellite images, leading to potential unresolved hallucination issues [62] in generated text descriptions. Besides, rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases. On the other hand, manual rewriting is highly labor-intensive.
@@
Therefore, an effective method for evaluating text quality and calibrating the alignment between text and image content is necessary to ensure a high level of LMM-inherent knowledge injection. In contrast to the common evaluation methods used to assess the quality of descriptions generated by LMMs, which typically rely on manually annotated image descriptions as references [65, 83, 34] to measure content overlap, we introduce an automatic mechanism designed to simulate human evaluation systems. This mechanism places greater emphasis on semantic similarity rather than merely assessing typical content overlap, thereby introducing PerceptionScore as a novel quality metric. Specifically, it comprises two components: text semantic quality and visual recall quality. The former approach directly adopts the efficient implementation of CLIPScore [25], which is based on CLIP’s [61] strong zero-shot capabilities and demonstrates high correlation with human semantic judgment. However, CLIPScore solely provides a measure of textual quality and fails to indicate whether specific elements in the image are omitted in the text (visual recall) [26].
@@
[…]
$\begin{split}&I^{\prime}=\operatorname{SDXL}(I),\quad\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),\\ &\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+\mathop{\texttt{CycleScore}(I,I^{\prime})})/2.\end{split}$ (1)
3.2 Multi-Granularity Cross-modal Alignment
@@
Modality Representation. Utilizing automated generation techniques, we obtain a dataset of high-quality image-text pairs, represented as $D=(I,T)$. Here, $I$ denotes satellite images $I^{st}$ or street-view images $I^{sv}$, while $T$ represents satellite text descriptions $T^{st}$ or street-view text descriptions $T^{sv}$.
@@
We first deploy Vision Transformer (ViT) [15] as an image encoder to process images derived from satellite and street-view sources. Finally, we use the output $\mathbf{z}_{l}^{0}$ corresponding to the special tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation. Concurrently, we use basic Transformer-Encoder [68] as a text encoder by default to comprehend and encode the semantics of textual descriptions in parallel. Furthermore, considering that the geospatial location information associated with street-view images can aid in pinpointing the precise geospatial coordinates of the area, we utilize the open-sourced GeoCLIP’s [69] location encoder to effectively instill geospatial information. This location encoder captures longitude and latitude $\mathcal{L}_{g}$, generating semantically rich, high-dimensional geospatial features, denoted as $\mathbf{z}_{L}$.
@@
Modality Alignment. The efficient alignment of visual and textual modalities is crucial for accurately modeling regional representations at both the coarse-grained global level and the fine-grained local level. To capture global visual information, in addition to the overall visual features for the region derived from satellite images, the clustering aggregation of location-level clues gathered within the region can additionally inject fine-grained details. Therefore, we further incorporate fine-grained information from street-view images, which are simultaneously fused with the corresponding geographical features. This approach aids in learning a robust and comprehensive representation of urban areas. The formalized process is as follows:
$\mathbf{z}_{g_{st}}=\textit{f}~{}(\mathbf{z}_{I_{st}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},\mathbf{z}_{I_{sv}}^{2},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},\mathbf{z}_{L}^{2},...,\mathbf{z}_{L}^{m})),$ (2)
@@
where $\mathbf{z}_{I_{st}}$ represents the satellite visual features for the queried region, $\mathbf{z}_{I_{sv}}^{k}$ denotes the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street-view location feature in the same region. The variable $m$ indicates the count of street-view scenes within the region. $\operatorname{Aggr}(\cdot)$ denotes the aggregation method. The function f  signifies the feature fusion method.
@@
To establish a global-level semantic correspondence between satellite visual data and their associated textual descriptions, we employ a joint contrastive optimization approach for the image and text modal encoders in the satellite branch. For the $i$-th satellite image-text pair $(I^{st}_{i},T^{st}_{i})$ in a mini-batch, we contrast the image-text pairs against others in the samples, which aims to maximally preserve the mutual information between the authentic pairs in the latent space. The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measure the similarity between the visual and textual embeddings, respectively, defined as: The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms which measure the similarity between the visual and textual embeddings, respectively, defined as:
$\mathcal{L}_{CG}=-\frac{1}{N}(\sum_{i}^{N}{\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}}}+\sum_{i}^{N}{\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}}}),$ (3)

### v3

# UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Socioeconomic Indicator Prediction

@@ Abstract
Urban socioeconomic indicator prediction aims to infer various metrics related to sustainable development in diverse urban landscapes using data-driven methods. However, prevalent pretrained models, particularly those reliant on satellite imagery, face dual challenges. Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place. Secondly, the text generated by the precursor work UrbanCLIP, which fully utilizes the extensive knowledge of LLMs, frequently exhibits issues such as hallucination and homogenization, resulting in a lack of reliable quality. In response to these issues, we devise a novel framework entitled UrbanVLP based on Vision-Language Pretraining. Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pretrained models. Moreover, it introduces automatic text generation and calibration, providing a robust guarantee for producing high-quality text descriptions of urban imagery. Rigorous experiments conducted across six socioeconomic indicator prediction tasks underscore its superior performance.
Codehttps://github.com/CityMind-Lab/UrbanVLP
Introduction
Urban Socioeconomic Indicator (USI) Prediction, a scholarly pursuit in the realm of urban computing (Zou et al. 2024), deploys data-driven methodologies to forecast socioeconomic metrics (e.g., GDP, population, carbon emission). Rooted in the escalating global prominence of urban environments and the imperative for judicious urban planning, this academic discipline finds motivation in enhancing the efficacy of policy-making, optimizing resource allocation, and mitigating challenges endemic to sustainable development (Xi et al. 2022; Yan et al. 2024).
@@ Introduction
Figure 1: USI prediction frameworks. Compared to existing arts, we present the first attempt to introduce multi-granular visual information and high-quality calibrated texts.
Given the widespread accessibility of satellite imagery on platforms such as Google Maps (Li et al. 2022c; Liu et al. 2023d; Xi et al. 2022), coupled with its wealth of information on regional features (e.g., road networks, building density, and vegetation coverage), a predominant majority of initiatives harness satellite imagery as the foundational modality for learning urban region representations and making predictions (Zhang et al. 2024). Through a comprehensive review of the current literature, we summarize mainstream methods for USI prediction into two categories:
  • Task-specific supervised learning commonly realizes USI prediction in a fully supervised task, e.g., identifying poverty levels (Han et al. 2020; Ayush et al. 2020, 2021), crop yields (Rußwurm and Körner 2020; Martinez et al. 2021; M Rustowicz et al. 2019; Yeh et al. 2021), and commercial activeness (He et al. 2018; Liu et al. 2023d). However, its inherent task specificity (i.e., reliant on ample labeled data), may impede the model’s capacity for broad generalization to other downstream tasks.
  • Vision pretraining, also known as a type of Self-Supervised Learning (Jaiswal et al. 2020), aims to learn general visual features from urban satellite imagery, which are subsequently fine-tuned on a specific task for enhancing performance  (Bai et al. 2023; Xi et al. 2022; Liu et al. 2023d). The inclusion of more auxiliary data, e.g., Points of Interests (POIs) (Zhang et al. 2021; Li et al. 2024b) and road networks (Liang et al. 2018, 2021; Chen et al. 2024), leads to richer, more accurate, and more useful representations of urban areas. For clarity, Figure 1 (a-b) depict a sketch of these two streams of approaches.
Despite the success of utilizing satellite imagery for USI prediction, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture). Previous frameworks in Figure 1 (a-b) mainly focus on single granularity, neglecting finer-grained visual clues. Nonetheless, as the saying goes, “the devil is in the details.”. By zooming into micro levels using corresponding street-view images, a more nuanced and fine-grained understanding emerges, as shown in the Figure 1(d) Left. Therefore, the integration and alignment of multi-granularity information remain ongoing challenges in USI prediction that require further exploration.
In recent years, by leveraging the extensive knowledge embedded in LLMs (Manvi et al. 2023) and the inherent interpretability of text modality, a wide range of tasks across various domains are empowered. Textual data can serve as an effective auxiliary tool in multimodal learning across a wide range of scenarios, such as geo-localization (Li et al. 2024a) and recommender systems (Gao et al. 2024).
While in USI prediciton, the efficacy of leveraging textual modality for semantic enrichment is still less explored. UrbanCLIP (Yan et al. 2024) stands out as the innovative solution, which generates text description by a pretrained LLM for satellite imagery and achieves learning urban region representations through natural language supervision.
@@
Nevertheless, the text generation process of UrbanCLIP raises several critical concerns: i) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery. ii) Homogenization: The generated descriptions appear overly simplified and general, potentially leading to homogenization and inadequate differentiation. To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
@@
In this paper, we present a Vision-Language Pretraining framework (i.e., UrbanVLP) for urban socioeconomic indicator prediction. As depicted in Figure 1(d), our model elaborately integrates multi-granularity information from both satellite (macro-level) and street-view (micro-level) imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions. Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module, which utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities. To address the second issue (i.e., hallucination and homogenization in LLM generated texts), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards. To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric called PerceptionScore, motivated by the human evaluation system.
In summary, our contributions lie in the following aspects:
@@
  • Multi-Granularity Cross-Modal Alignment. We explore the role of two distinct visual data modalities at various semantic granularities – satellite imagery and street-view imagery. We inject fine-grained semantic information by integrating street-view data through token-level contrastive learning.
@@
  • Automatic Text Generation and Calibration. Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism based on a new reference-free metric, ensuring the fidelity between the generated texts and the corresponding image content.
  • A New Benchmark & Empirical Evidence. We plan to open-source the first vision-text and multi-granularity urban dataset upon paper notification, consisting of six downstream tasks across the socio-economy. Extensive experiments demonstrate that our UrbanVLP outperforms existing approaches by an average improvement of 3.95% on the $R^{2}$ metric, as illustrated in Figure 2.
Figure 2: $R^{2}$ results in Beijing and Shenzhen.
@@
Related Work
We formally define the problem of USI prediction. Given a satellite image $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$. The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$. The representation can then be further utilized in downstream tasks, such as inferring socioeconomic indicators $\mathbf{Y}_{g}$ for the given region $g$. The detailed formulation and definitions related to this task are provided in Appendix A.
Urban Socioeconomic Indicator Prediction. In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Liu et al. 2023d; Xi et al. 2022). For example, Urban2Vec (Wang, Li, and Rajagopal 2020) integrated street-view imagery and POI data to learn neighborhood embeddings. Some contrastive learning approaches like PG-SimCLR (Xi et al. 2022) and UrbanCLIP (Yan et al. 2024) have shown success in representing urban regions through satellite images. However, the aforementioned works exclusively considered a single type of urban imagery, overlooking the potential synergy between satellite imagery and street-view imagery, which can complement each other.
Vision-Language Pretraining (VLP). VLP aims to jointly encode vision and language in a fusion model. A milestone work CLIP (Radford et al. 2021) and its variants (Li et al. 2021; Yao et al. 2022) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval. Recent works (Tsimpoukelli et al. 2021; Alayrac et al. 2022) shift towards leveraging LLMs knowledge for VLP. In USI prediction, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information.
Methodology
@@
Figure 3 shows our framework with two stages. In the pretraining stage, we first devise an automatic text generation and calibration module using ShareGPT4V (Chen et al. 2023b) to generate textual descriptions for street-view images with geographical and visual prompts. To guarantee the quality of the generated text, we introduced a novel metric called PerceptionScore. Then a multi-granularity cross-modal alignment framework is designed to utilize multi-level contrastive learning and fine-grained information injection. During the fine-tuning stage, we employ frozen encoders from Stage 1 to extract features and fine-tune a lightweight MLP for accurate predictions.
Automatic Text Generation and Calibration
@@
Text Generation. For each street-view image, we use advanced Large Multimodal Models (LMMs) due to their robust cross-modal knowledge retention abilities, to provide comprehensive textual descriptions. However, this confronts two primary challenges. Firstly, due to substantial financial costs associated with API usage, employing closed source LMMs like GPT4V (Yang et al. 2023) or Gemini (Fu et al. 2023) for generating tens of thousands of text descriptions is impractical. Secondly, street-view images often encompass diverse details, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
@@
To address these issues, we first employ an open-source model closely related to GPT-4V, known as ShareGPT4V (Chen et al. 2023b), to generate the street-view image descriptions. This model is a powerful LLM with vision capabilities trained on 100k GPT4V-generated captions, revealing captioning ability comparable to GPT4V. Moreover, in the process of designing text prompt templates, we believe that the proportion of each element in street-view images reflects their relative abundance, thereby facilitating LMMs in evaluating their significance. Therefore, we employ a pretrained segmentation model (CSAILVision 2024) to decouple visual elements and calculate the proportion of segmentation for each element. As evidenced by existing research (Fan et al. 2023b), the accuracy of current segmentation methods is already adequate to recognize various visual components. Simultaneously, geospatial coordinates of street-view images are incorporated as prompts to aid in the generation of more precise textual descriptions. Template and segmentation ratio details can be found in Appendix H.
@@
Text Calibration. Existing works (Fan et al. 2023a; Yan et al. 2024) have already demonstrated that the quality of the textural descriptions is crucial for enhancing model performance. However, previous research (Yan et al. 2024) primarily employs simplistic rules and manual refinement processes for refining text description of satellite imagery, leading to potential unresolved hallucination problems (Rawte, Sheth, and Das 2023). Besides, the rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases. On the other hand, manual rewriting is highly labor-intensive, which limits its scalability.
@@
Therefore, an effective method for assessing the quality of generated text and calibrating the alignment between the text and image is essential to ensure a high level of LMM-inherent knowledge integration. Compared to classical methods that assess LMM-generated descriptions by relying on manually-annotated references to measure content overlap (Zhang et al. 2020; Lee et al. 2020), we devise an automatic mechanism to simulate human evaluation systems. This mechanism places greater emphasis on semantic consistency rather than merely assessing typical content overlap, by introducing a novel quality metric called PerceptionScore. Concretely, it comprises two parts: text semantic quality and visual recall quality. The former one directly adopts the efficient implementation of CLIPScore (Hessel et al. 2021), which is based on CLIP’s strong zero-shot capabilities and demonstrates high correlation with human semantic judgment. However, CLIPScore solely measures textual quality without consideration of the omission of specific visual elements in the text, thereby lacking an assessment of visual recall (Hu et al. 2023).
[…]
$I^{\prime}=\operatorname{SDXL}(I),$ (1)
$\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),$ (2)
$\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+$
$\mathop{\texttt{CycleScore}}(I,I^{\prime}))/2.$ (3)
@@
Multi-Granularity Cross-modal Alignment
@@
Modality Representation. Utilizing automated text generation methods, we obtain a dataset of high-quality image-text pairs $D=(I,T)$. Here, $I$ represents satellite images $I^{st}$ or street-view images $I^{sv}$, while $T$ indicates satellite text descriptions $T^{st}$ or street-view text descriptions $T^{sv}$.
@@
We first deploy Vision Transformer (ViT) (Dosovitskiy et al. 2021) as the image encoder to process images derived from satellite and street-view sources. Then we use the output $\mathbf{z}_{l}^{0}$ corresponding to the global tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation. Concurrently, we leverage the basic Transformer-Encoder (Vaswani et al. 2017) as the text encoder to comprehend and encode the semantics of textual descriptions in parallel. Furthermore, considering that the geospatial location information associated with street-view images can assist in pinpointing the precise geospatial coordinates of the area, we utilize the open-sourced GeoCLIP’s (Vivanco, Nayak, and Shah 2023) location encoder to capture longitude and latitude $\mathcal{L}_{g}$, generating semantically informative geospatial features, denoted as $\mathbf{z}_{L}$. Additional details can be found in Appendix G.
@@
Modality Alignment. To capture the comprehensive visual information, beyond deriving global visual features for the region from satellite images, the aggregation of location-level clues within the region can enhance the injection of details. Therefore, we further incorporate fine-grained information from street-view images, along with the corresponding geographical features. The formalized process is as follows:
$\mathbf{z}_{g_{st}}=\textit{f}~{}(\mathbf{z}_{I_{st}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},...,\mathbf{z}_{L}^{m})),$ (4)
@@
where $\mathbf{z}_{I_{st}}$ denotes the satellite visual features for the query region, $\mathbf{z}_{I_{sv}}^{k}$ is the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street-view location feature in the same region. $m$ indicates the count of street-view scenes within the region. $\operatorname{Aggr}(\cdot)$ denotes the aggregation method. The function f  signifies the feature fusion method.
@@
To establish a global-level semantic correspondence between the satellite features and their associated textual features, we employ a joint contrastive optimization approach for the image and text encoders in the satellite branch. For the $i$-th satellite image-text pair $(I^{st}_{i},T^{st}_{i})$ in a mini-batch, we contrast the image-text pairs against others within the samples, which aims to maximally preserve the mutual information between the pairs in latent space. The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measures the similarity between the visual and textual embeddings, respectively, defined as:
$\mathcal{L}_{CG}=-\frac{1}{N}\Bigg{(}$ $\sum_{i}^{N}\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}$ (5)
$+\sum_{i}^{N}\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}\Bigg{)},$